# Precise Constraints on the Energy Budget of WASP-121 b from its JWST NIRISS/SOSS Phase Curve

### v1

@@ Acknowledgments
This project was undertaken with the financial support of the Canadian Space Agency through a NEAT/GTO grant. The contributions from RF were supported in part by Grant #2019-1403 from the Heising-Simons Foundation. N.B.C. acknowledges support from an NSERC Discovery Grant, a Tier 2 Canada Research Chair, and an Arthur B. McDonald Fellowship and thanks the Trottier Space Institute and l’Institut de recherche sur les exoplanètes for their financial support and dynamic intellectual environment. L.D. is a Banting and Trottier Postdoctoral Fellow and acknowledges support from the Natural Sciences and Engineering Research Council (NSERC) and the Trottier Family Foundation. S.P. acknowledges support from the Swiss National Science Foundation under grant 51NF40_205606 within the framework of the National Centre of Competence in Research PlanetS. R.J.M. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51513.001, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS 5-26555. D.L acknowledges financial support from NSERC and FRQNT. C.P.-G. acknowledges support from the E. Margaret Burbidge Prize Postdoctoral Fellowship from the Brinson Foundation. JDT acknowledges funding support by the TESS Guest Investigator Program G06165. Some/all of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/ppjp-9743 (catalog doi: 10.17909/ppjp-9743).
@@ Appendix A NAMELESS data reduction and light-curve fitting
An additional reduction of the data is produced using the NAMELESS data reduction pipeline (Coulombe et al., 2023, 2025) following the methodology described in Allart et al. (Under Review). In short, we go through all Stage 1 and Stage 2 steps of the STScI jwst pipeline until the flatfield correction step. We then correct for bad pixels by interpolating over pixels whose mixed spatial second derivative are away from the local median. The non-uniform background is subtracted from all integrations by scaling uniformly the two portions of the model background provided by STScI that are separated by the sharp jump in background flux near . Cosmic rays are corrected for by computing the running median of all pixels and bringing any count that is a outlier to the value of its pixel’s running median. We correct 1/ noise by scaling individually each column of the order 1 and 2 traces and computing the constant additive that best reduces the chi-square (considering only pixels that are pixels away from the center of the trace), which is then subtracted from the data. Finally, we extract the stellar spectra using a fixed box aperture of width 40 pixels. The order 1 and 2 white-light curves produced from this reduction are shown in Figure A3.
Two sets of spectroscopic light curves are produced by binning the data at fixed resolving powers of and . The full fitting methodology is presented in Pelletier et al. (2025b) and is described briefly below. The astrophysical model consists of a second order sinusoid (, where is the orbital phase) which is combined to a transit and eclipse model computed using batman (Kreidberg, 2015). We adopt large uniform priors for the coefficients of the sinusoidal model, allowing for negative flux. The transit parameters consist of the planet-to-star radius ratio () and the quadratic limb-darkening coefficients [, ] (, following Coulombe et al. (2024)). For the systematics model, we consider a linear slope in time and a Heaviside step function centered at the time where the tilt event occurs (). We also freely fit the photometric scatter for each spectral bin. The spectroscopic fits are ran assuming zero-eccentricity (), a mid-transit time of MJD, an orbital period of days (Patel & Espinoza, 2022), as well as values of and for the scaled semi-major axis and impact parameter as measured from the NAMELESS order 1 and 2 white-light curves. The light-curve fits are ran with the emcee Python package (Foreman-Mackey et al., 2013) for 6,250 steps considering 4 walkers per free parameter. All free parameters are fit considering wide uniform priors. The resulting day- and nightside spectra for the fit are shown in Figure A5 and the phase curve offset spectrum from the fit is shown in Figure A7.
@@ Appendix B Analytical Solutions to Thermal Emission and Reflected Light Kernels
Here, we present analytical solutions to the thermal emission and reflected light kernels presented in section 2.4 in equations 12 & 13.
@@ B.2 Reflected Light Kernel
This solution also carries the same caveats that the thermal emission kernel solution had to remain physical. Namely, the slice edges must be within the visibile hemisphere, .

### v2

@@ Acknowledgments
This project was undertaken with the financial support of the Canadian Space Agency through a NEAT/GTO grant. The contributions from RF were supported in part by Grant #2019-1403 from the Heising-Simons Foundation. N.B.C. acknowledges support from an NSERC Discovery Grant, a Tier 2 Canada Research Chair, and an Arthur B. McDonald Fellowship and thanks the Trottier Space Institute and l’Institut de recherche sur les exoplanètes for their financial support and dynamic intellectual environment. L.D. is a Banting and Trottier Postdoctoral Fellow and acknowledges support from the Natural Sciences and Engineering Research Council (NSERC) and the Trottier Family Foundation. S.P. acknowledges support from the Swiss National Science Foundation under grant 51NF40_205606 within the framework of the National Centre of Competence in Research PlanetS. R.J.M. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51513.001, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS 5-26555. R.A. acknowledges the Swiss National Science Foundation (SNSF) support under the Post-Doc Mobility grant P500PT_222212 and the support of the Institut Trottier de Recherche sur les Exoplanètes (IREx). D.L acknowledges financial support from NSERC and FRQNT. C.P.-G. acknowledges support from the E. Margaret Burbidge Prize Postdoctoral Fellowship from the Brinson Foundation. JDT acknowledges funding support by the TESS Guest Investigator Program G06165. Some/all of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/ppjp-9743 (catalog doi: 10.17909/ppjp-9743).
@@ Appendix A NAMELESS data reduction and light-curve fitting
An additional reduction of the data is produced using the NAMELESS data reduction pipeline (Coulombe2023; Coulombe2025-LTT-NIRISS-PC) following the methodology described in Allart2025. In short, we go through all Stage 1 and Stage 2 steps of the STScI jwst pipeline until the flatfield correction step. We then correct for bad pixels by interpolating over pixels whose mixed spatial second derivative are away from the local median. The non-uniform background is subtracted from all integrations by scaling uniformly the two portions of the model background provided by STScI that are separated by the sharp jump in background flux near . Cosmic rays are corrected for by computing the running median of all pixels and bringing any count that is a outlier to the value of its pixel’s running median. We correct 1/ noise by scaling individually each column of the order 1 and 2 traces and computing the constant additive that best reduces the chi-square (considering only pixels that are pixels away from the center of the trace), which is then subtracted from the data. Finally, we extract the stellar spectra using a fixed box aperture of width 40 pixels. The order 1 and 2 white light-curves produced from this reduction are shown in Figure C3.
Two sets of spectroscopic light-curves are produced by binning the data at fixed resolving powers of and . The full fitting methodology is presented in Pelletier2025-NIRISS-emission and is described briefly below. The astrophysical model consists of a second order sinusoid (, where is the orbital phase) which is combined to a transit and eclipse model computed using batman (batman). We adopt large uniform priors for the coefficients of the sinusoidal model, allowing for negative flux. The transit parameters consist of the planet-to-star radius ratio () and the quadratic limb-darkening coefficients [, ] (, following Coulombe2024). For the systematics model, we consider a linear slope in time and a Heaviside step function centered at the time where the tilt event occurs (). We also freely fit the photometric scatter for each spectral bin. The spectroscopic fits are ran assuming zero-eccentricity (), a mid-transit time of MJD, an orbital period of days (Patel2022), as well as values of and for the scaled semi-major axis and impact parameter as measured from the NAMELESS order 1 and 2 white-light-curves. The light-curve fits are ran with the emcee Python package (emcee) for 6250 steps considering 4 walkers per free parameter. All free parameters are fit considering wide uniform priors. The resulting day- and nightside spectra for the fit are shown in Figure C5 and the phase curve offset spectrum from the fit is shown in Figure C7.
@@ Appendix B Analytical Solutions to Thermal Emission and Reflected Light Kernels
Here, we present analytical solutions to the thermal emission and reflected light kernels presented in section 2.4 in equations 12 and 13.
@@ B.2 Reflected Light Kernel
This solution also carries the same caveats that the thermal emission kernel solution had to remain physical. Namely, the slice edges must be within the visible hemisphere, .
@@ Appendix C Additional Plots
Figures C1–C9 provide additional visualizations and checks that complement the results discussed in the main sections.